OpenStudy (anonymous):

solve 0.011 ^-2, the answer is not important, I just want to know now to solve these equations.

OpenStudy (anonymous):

the first thing you have to know is what a negative exponent means. \[b^2\] is pretty straightforward: is it \[b^2=b\times b\] but the minus sign in the exponent means to take the reciprocal, so \[b^{-2}=\frac{1}{b^2}\]

OpenStudy (anonymous):

in this example \[b=0.011\] so if you want \[(0.011)^{-2}\] you have to compute \[\frac{1}{(0.011)^2}\]

OpenStudy (anonymous):

hope that helped

OpenStudy (anonymous):

erm I could still do with some help

OpenStudy (anonymous):

so 0.011^2= 0.000121

OpenStudy (anonymous):

yes but you did have an exponent of -2 right? not 2

OpenStudy (anonymous):

so it is 1/0.000121?

OpenStudy (anonymous):

Opps.... I missed the - sign in the exponent...

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

it is \[\frac{1}{.000121}\] and this you can compute by using a calculator, or by changing it to a fraction as \[\frac{1,000,000}{121}\]

OpenStudy (anonymous):

[1/(0.011)^2] =[1/(11/1000)^2]

OpenStudy (anonymous):

oh that is a fraction, I thought it meant divide

OpenStudy (anonymous):

@hollywood they mean the same thing

OpenStudy (anonymous):

\[\frac{5}{2}\] means "five halves" or "five divided by two"

OpenStudy (anonymous):

a fraction bar is a synonym for division. \[2\div 3=\frac{2}{3}\]

OpenStudy (anonymous):

so \[\frac{1}{0.000121}=1\div0.000121\]

OpenStudy (anonymous):

if you wanted to compute this using a calculator you would press the \[\div\] button

OpenStudy (anonymous):

in the last problem i posted, the 1 became 10000000

OpenStudy (anonymous):

ok that is because if i want to write \[\frac{1}{.02}\] for example as a fraction, i would move the decimal place two spaces right in the numerator and denominator

OpenStudy (anonymous):

you might think "why isn't \[\frac{1}{.02}\] a fraction already?" that is because you have a fraction and a decimal

OpenStudy (anonymous):

so moving over two spaces i write \[\frac{1}{.02}=\frac{100}{2}\] now i have a fraction

OpenStudy (anonymous):

and now i see that since \[\frac{100}{2}=100\div 2=50\] i know what that number is

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

so you divide 121 by 1000000